# Autoregressive Action Sequence Learning for Robotic Manipulation

### v2

@@ Abstract
Autoregressive models have demonstrated remarkable success in natural language processing. In this work, we design a simple yet effective autoregressive architecture for robotic manipulation tasks. We propose the Chunking Causal Transformer (CCT), which extends the next-single-token prediction of causal transformers to support multi-token prediction in a single pass. Further, we design a novel attention interleaving strategy that allows CCT to be trained efficiently with teacher-forcing. Based on CCT, we propose the Autoregressive Policy (ARP) model, which learns to generate action sequences autoregressively. We find that action sequence learning enables better leverage of the underlying causal relationships in robotic tasks. We evaluate ARP across diverse robotic manipulation environments, including Push-T, ALOHA, and RLBench, and show that it outperforms the state-of-the-art methods in all tested environments, while being more efficient in computation and parameter sizes. Video demonstrations, our source code, and the models of ARP can be found at http://github.com/mlzxy/arp.
@@ 1 Introduction
Autoregressive models are the basis of recent breakthroughs in natural language processing (Min et al., 2023). These models predict the next token in a given sequence based on the previous tokens. Autoregressive models are implemented with causal transformers, where each token attends only to preceding ones, and they are trained with the single objective of maximizing the conditional likelihood of each token. Despite their simplicity, autoregressive models such as GPTs (Mann et al., 2020) are shown to demonstrate a reasoning ability that can capture causal dependencies (Prystawski et al., 2024). In this work, we explore the design of a simple autoregressive architecture that can be used for various robot manipulation tasks in diverse environments.
Decision Transformer (DT) and Trajectory Transformer (TT) are the pioneering approaches that use autoregressive models to solve control tasks (Chen et al., 2021; Janner et al., 2021). These methods learn to generate trajectories as $(R_{1},s_{1},a_{1},\dots,R_{T},s_{T},a_{T})$, where $R_{t},s_{t},a_{t}$ respectively denote the reward-to-go (Tamar et al., 2016), the state, and the action at time-step $t$. In contrast, we propose to predict only the future action sequence, and condition the prediction on the current state (or observation). Action sequence learning is more attainable as an objective in robotics applications, where the underlying reward functions are unknown or ill-defined, and observations, given as images or point clouds, are high-dimensional and cannot be predicted accurately (Kroemer et al., 2021). DT and TT are mainly used in tasks where low-dimensional state variables are fully observed. To deal with uncertainty, our model generates a new action sequence after every $k$ time-steps, using an updated new observation and following the Model Predictive Control (MPC) approach. Action sequence modeling also enables a better leverage of the underlying causal dependencies in robotic tasks. Examples of such causal dependencies include: logical dependencies, where low-level actions depend on high-level plans, spatial dependencies, where orientation depends on the end-effector’s position, and temporal dependencies, where latter actions depend on earlier ones. We showcase the action sequence designs of our real robot experiment, Push-T, ALOHA, and RLBench in Figure 3.
We propose the Chunking Causal Transformer (CCT), an auto-regressive model that is tailored for robotic tasks. CCT extends the next-token predictions of causal transformer to support multi-token predictions in a single pass. CCT predicts the future tokens, or a chunk of actions, from empty tokens rather than from the original sequence, as illustrated in Figure A1. In doing so, CCT is more efficient by requiring fewer inference passes and delivers stronger performance because grouping actions in chunks is critical for robotic tasks that require high control frequency (Zhao et al., 2023). Further, we design a novel attention interleaving strategy that allows CCT to be trained efficiently with teacher-forcing, as shown in Figure 4. While action chunking has been previously introduced in the Action Chunking Transformer (ACT) by Zhao et al. (2023), ACT is a one-step prediction model. Instead, we employ action chunking for auto-regressive models. Our ablation studies show that both auto-regression and action chunking are the key factors behind the strong performance of the proposed model, as shown in Table 1 and Figure 7.
@@
To summarize, our contributions are threefold. (1) We propose the Chunking Causal Transformer (CCT), which extends the next-token prediction of causal transformer to multi-token prediction for auto-regressive models. We also design a novel attention interleaving strategy that allows CCT to be trained efficiently with teacher-forcing. (2) Based on our CCT, we present the Auto-regressive Policy (ARP), a model that learns to generate action sequences auto-regressively for solving robotic manipulation tasks. The ARP architecture is summarized in Figure 1. (3) We evaluate ARP across Push-T (Chi et al., 2023), ALOHA (Zhao et al., 2023), and RLBench (James et al., 2020), three environments with diverse manipulation tasks, as outlined in Figure 2. Our study shows that ARP outperforms all environment-specific SoTAs, while being more efficient computationally and using smaller parameter sizes, as summarized in Figure 5. In addition, we evaluate ARP with a real robot on a challenging, contact-rich nut-tightening task, as shown in Figure 8.
Figure 1: Autoregressive Policy Architecture. A sequence of past actions and a chunk of empty tokens are concatenated and projected into embeddings. The empty tokens correspond to future actions, which are unknown and need to be predicted. These embeddings are fed into our Chunking Causal Transformer (CCT) along with the vision features of the current observation. CCT alternates between self-attention within the input embeddings and cross-attention with the vision features. Self-attention is causal for the known input actions and bidirectional among the empty tokens. Distributions of future actions are decoded from the updated embeddings of the empty tokens.
@@ 2 Related Work
Learning robotic manipulation from demonstrations. Imitation learning enables robots to learn to perform tasks demonstrated by experts (Zare et al., 2024). Recently, various methods have been developed for manipulation learning with different task constraints and control modalities. Notably, Chi et al. (2023) proposed the Diffusion Policy (DP) method for solving the Push-T task. Zhao et al. (2023) proposed the Action Chunking Transformer (ACT) for bi-manual manipulation tasks in the ALOHA environment. Goyal et al. (2024) proposed RVT-2 for language-conditioned tasks in the RLBench environment (James et al., 2020). We outline these environments and the corresponding state-of-the-art (SoTA) solutions in Figure 2 and Figure 6, respectively. In contrast, our proposed auto-regressive policy is a universal architecture that outperforms each environment-specific SoTA on Push-T, ALOHA, and RLBench.
Sequence models for control tasks. In addition to the Decision Transformer (DT) and the Trajectory Transformer (TT), recent works such as OpenVLA (Kim et al., 2024) and ManipLLM (Li et al., 2024) have looked into fine-tuning a large language model (LLM) such as LLaMA (Li et al., 2024) to directly include target end-effector poses within text-based responses. Despite their impressive results, these approaches are limited to low-frequency control tasks that rely on end-effector waypoints (Kim et al., 2024). Moreover, the reliance on resource-intensive LLMs leads to large computational overhead, even for tasks that could be addressed with lightweight models. Without these constraints, our auto-regressive policy outperforms SoTAs in multiple environments while being more efficient in MACs (number of multiply-accumulate operations) and parameter sizes.
Hierarchical policies. Planning actions on multiple levels of abstraction is an important ability (Pateria et al., 2021). Existing methods generally separate the designs of low-level and high-level policies, and uses different modules for the different levels of abstraction (Le et al., 2018; Pateria et al., 2021; Belkhale et al., 2023; 2024; Chen et al., 2024b). This complicated procedure prohibits a wider application. In contrast, our auto-regressive policy predicts a sequence of actions of different levels of abstraction by using a single model.
@@ 3 Method
Action sequence design. Unlike natural language, robot actions lack a universal vocabulary. As shown in Figure 2, different robot tasks may require drastically different types of actions. We therefore propose to represent actions as structured sequences that follow a format that is pre-defined for each family of tasks. Figure 3 showcases the formats of the action sequences generated in our real robot experiment, Push-T, ALOHA, and RLBench tasks.
Action embedding and decoding. Language models map each word to a continuous vector called word embedding. The word embeddings of the input sentences are fed into a causal transformer. The distribution of the next word is decoded from the output embedding of the last word with a linear layer. Figure A2 and A3 illustrate our embedding and decoding methods for robot actions. Discrete actions are embedded by a table lookup on a weight matrix and decoded into a categorical distribution with a linear layer, similar to words in language modeling. Continuous actions are embedded with a linear layer and decoded into the parameters of a Gaussian mixture distribution with another linear layer. Actions that are defined as pixel coordinates are embedded by retrieving the point-wise features at the coordinates on a visual feature map. The output spatial distribution is obtained by multiplying the output embedding with the visual feature map, and converting the result into a 2-d heatmap with the up-sampling operator from RAFT (Teed & Deng, 2020).
@@
Chunking causal transformer. Figure A1 illustrates the essential difference between a causal transformer and our CCT. A causal transformer modifies the token embedding with causal attention so that the last token becomes the next token. Our CCT modifies the token embedding with causal attention for the known tokens $a_{i}$ (past actions) and bidirectional attention for the empty tokens $e_{i}$ (future actions). The empty tokens become the next tokens. This allows the prediction of multiple next tokens at once in a single forward pass by adding empty tokens. The advantages are two-fold: (1) A better accuracy is achieved because error accumulation is reduced when actions are grouped in chunks and executed as one unit (Zhao et al., 2023). (2) A better efficiency is achieved with fewer forward passes because each forward pass predicts multiple tokens at once. We study the impacts of action chunking in detail in Section 4. In ARP, CCT alternates between self-attention within the input embeddings and cross-attention with vision features, as in Figure 1. We extract vision features from a standard backbone identical to the ones used in SoTA methods, as detailed in section 4.
@@
To resolve this, we first divide the attention into three categories: (1) causal attention within the known tokens, (2) bidirectional attention within the empty tokens, and (3) causal attention between the empty and the known tokens, as marked in Figure 4. Figure A4 shows that the causal attention within known tokens is computed repeatedly. We avoid this redundancy by precomputing the causal attentions for all known tokens and caching the results. For the empty tokens, we combine their inner bidirectional attention and the causal attention toward cached known tokens. This enables a single forward pass of all tokens in three attention operations, regardless of the number of tokens. We name this procedure attention interleaving. Figure A4 demonstrates the reduced MACs of training with attention interleaving. We implement attention interleaving as an internal acceleration mechanism of the transformer layer, which is transparent to other network modules. Note that attention interleaving is only used during training and incurs no additional inference cost.
Inference. During the test rollouts, we extract vision tokens from the current observation and provide them as input to ARP, which then generates actions auto-regressively by sampling from the decoded action distribution and appending the selected actions to the existing action sequence. This process of generating and executing actions is repeated until episode termination (success, failure, or reaching the time limit). Actions are generated according to the sequence formats shown in Figure 3 with pre-defined chunk sizes and target lengths per task. We provide additional implementation details and hyper-parameter values in Section 4 and Appendix A.2.
Figure 4: Training Chunking Causal Transformer (CCT) with Teacher-forcing. Causal transformers are trained efficiently with only a single forward pass for all tokens in a given sequence. However, suppose $a_{2},a_{3}$ and $a_{4}$ are in separate chunks, the CCT forward passes of predicting $a_{2},a_{3}$ and $a_{4}$ cannot be merged directly. Naively running separate passes significantly increases computation costs, as in Figure A4. With the proposed attention interleaving, we precompute and cache the causal attention results for all known tokens $a_{1},a_{2},a_{3}$. For the empty tokens $e_{2},e_{3},e_{4}$, we combine their bidirectional attention (lightgray) and the causal attention towards known tokens (green, blue). Since all the computations related to the known tokens are cached, we can update all the empty tokens in a single pass, regardless of the number of tokens. An example of attention interleaving is provided in the Video/attention-interleaving-tour.mp4 in the supplementary.
@@ 4.1 Comparison with State-of-the-Art
Setup. We compare the auto-regressive policy (ARP) against the SoTA solutions in Push-T, ALOHA, and RLBench environments. Push-T is a single task. ALOHA consists of two tasks: insertion and cube transfer. RLBench includes 18 tasks, each with multiple language variants. These environments are illustrated in Figure 2 and Figure A5. For Push-T and ALOHA, we train a separate policy for each task. For RLBench, a single policy is trained for all 18 tasks. In Push-T, the policy observes the last two $96\times 96$ RGB frames, and predicts a window of future 2-d pointer positions as actions. In ALOHA, the policy observes the current $480\times 640$ RGB frame and predicts a window of future 14-dimensional joint positions. In RLBench, the policy observes four RGBD $128\times 128$ images and predicts the next target end-effector pose and gripper states. Existing SoTA techniques in these environments are outlined in Figure 6. We use the same vision backbones as the SoTA solutions to extract vision tokens, namely ResNet50 (He et al., 2016) for Push-T and ALOHA, and Multi-View Transformer (Goyal et al., 2023) for RLBench. We use the same training data, number of episodes, optimizer configuration, and evaluation frequency as the SoTA solutions. We detail the full list of hyper-parameters, such as the number of layers, sequence lengths, chunk sizes, and optimizer setups in Appendix A.2. Success rates for Push-T and RLBench are averaged over three independent runs. ALOHA’s results are averaged over five runs.
Results. Figure 5 shows that our auto-regressive policy (ARP) outperforms environment-specific SoTAs while being more computationally efficient. Table A4 compares the per-task success rates of our ARP and RVT-2 (Goyal et al., 2024). In addition, we report the result of ARP+, which shares the same network definition with ARP but has more layers. The MACs / parameter sizes of RVT-2, ARP, ARP+ are 72.1M/57.1G, 71.9M/56.2G, and 73.8M/57.4G, respectively. Note that the originally reported success rate of RVT-2 was 81.6%. This was achieved by using a logical time-step as an input. Many existing works have followed this convention (Shridhar et al., 2023; Goyal et al., 2023). We train all our RLBench models without the input time-step. Thus, our RLBench models are more robust to variations of task initialization in real applications.

### v3

@@ Abstract
Designing a universal policy architecture that performs well across diverse robots and task configurations remains a key challenge. In this work, we address this by representing robot actions as sequential data and generating actions through autoregressive sequence modeling. Existing autoregressive architectures generate end-effector waypoints sequentially as word tokens in language modeling, which are limited to low-frequency control tasks. Unlike language, robot actions are heterogeneous and often include continuous values—such as joint positions, 2D pixel coordinates, and end-effector poses—which are not easily suited for language-based modeling. Based on this insight, we introduce a straightforward enhancement: we extend causal transformers’ single-token prediction to support predicting a variable number of tokens in a single step through our Chunking Causal Transformer (CCT). This enhancement enables robust performance across diverse tasks of various control frequencies, greater efficiency by having fewer autoregression steps, and lead to a hybrid action sequence design by mixing different types of actions and using a different chunk size for each action type. Based on CCT, we propose the Autoregressive Policy (ARP) architecture, which solves manipulation tasks by generating hybrid action sequences. We evaluate ARP across diverse robotic manipulation environments, including Push-T, ALOHA, and RLBench, and show that ARP, as a universal architecture, outperforms the environment-specific state-of-the-art in all tested benchmarks, while being more efficient in computation and parameter sizes. Videos of our real robot demonstrations, all source code and the pretrained models of ARP can be found at http://github.com/mlzxy/arp.
@@ 1 Introduction
Autoregressive models are the basis of recent breakthroughs in natural language processing (Min et al., 2023). These models predict the next token in a sequence based on the previous tokens. Autoregressive models are implemented with causal transformers, where each token attends only to preceding ones, and they are trained with the single objective of maximizing the conditional likelihood of each token. Despite their simplicity, autoregressive models such as GPTs (Mann et al., 2020) are shown to demonstrate a reasoning ability that can capture causal dependencies (Prystawski et al., 2024). In this work, we explore the design of a universal autoregressive architecture that can be used for various robot manipulation tasks in diverse environments.
Decision Transformer (DT) and Trajectory Transformer (TT) are the pioneering approaches that use autoregressive models to solve control tasks (Chen et al., 2021; Janner et al., 2021). These methods learn to generate trajectories as $(R_{1},s_{1},a_{1},\dots,R_{T},s_{T},a_{T})$, where $R_{t},s_{t},a_{t}$ respectively denote the reward-to-go (Tamar et al., 2016), the state, and the action at time-step $t$. However, these methods are primarily applied to tasks with fully observed, low-dimensional states—a condition rarely met in robotics applications. Recent work focuses on applying autoregression only on action sequences, such as Gato (Reed et al., 2022), VIMA (Jiang et al., 2022), ManipLLM (Li et al., 2024). Despite the impressive results, these methods remain limited to low-frequency control tasks, as they represent robot actions with key end-effector waypoints and generate one action at a time, similar to word generation in language modeling.
However, unlike language, robot actions are heterogeneous and include continuous values—such as joint positions, 2D pixel coordinates, and effector poses. Additionally, in high-frequency control tasks, continuous actions are expected to maintain temporal smoothness—a requirement absent in language modeling. To adapt autoregressive models for robotic tasks, we propose the Chunking Causal Transformer (CCT), an improved version of the causal transformer used in standard autoregressive models. CCT introduces a simple modification: it predicts the future tokens, or a chunk of actions, from empty tokens rather than from the original sequence, as illustrated in Figure 5. In doing so, CCT extends the next-single-token prediction of causal transformer to chunking autoregression—next-multi-token prediction in a single step. Despite the simplicity of our enhancement, chunking autoregression offers three key advantages:
  1. 1.
    Predicting multiple temporally correlated actions in a single step addresses the primary limitation of autoregressive models in high-frequency control tasks.
  2. 2.
    Chunking autoregression increases efficiency by reducing the inference passes required.
  3. 3.
    Variable chunk sizes enable a flexible action sequence design, such as mixing different types of actions and using a different chunk size for each type. For example, high-level actions, like sparse 2D waypoints, can be predicted first sequentially to guide the prediction of low-level actions, such as joint positions, in larger chunks. We show action sequence designs of our real robot task, Push-T, ALOHA, and RLBench in Figure 4.
While action chunking has been previously introduced in the Action Chunking Transformer (ACT) by Zhao et al. (2023), ACT is a one-step prediction model with a fixed chunk size. Instead, our approach supports variable chunk sizes for more powerful autoregressive models in generating hybrid action sequences. Table 2 shows that our method outperforms ACT by a significant margin in all environments. Table 1 shows that our chunking autoregression is the key factor behind the strong performance of the proposed model. We illustrate the essential difference between action chunking, standard autoregression, and our chunking autoregression in Figure 1.
Figure 1: Existing Works versus Our Autoregressive Policy. Action chunking models (left) predict all action tokens in a single step (Zhao et al., 2023; Kim et al., 2024; Wu et al., 2023). Standard autoregression models (middle) generate one action token in each step, which is inefficient and unsuitable for high-frequency control tasks Chen et al. (2021); Janner et al. (2021); Reed et al. (2022); Jiang et al. (2022); Li et al. (2024). Our proposed chunking autoregression (right) generates a chunk of variable number of action tokens per step, offering greater efficiency, strong performance across diverse tasks, and flexibility in designing hybrid action sequences. We compare the performance of these three action prediction strategies in Table 1. Note all models use Model Predictive Control to predict $L$ actions, execute them, update the observation, and then predict actions again. Autoregressive generation is performed without executing actions or changing the current observation.
@@
To summarize, our contributions are threefold. (1) We propose the Chunking Causal Transformer (CCT), which extends the single-token prediction of causal transformer to multi-token prediction, and therefore enables chunking autoregression. We also design a novel attention interleaving strategy that allows CCT to be trained efficiently with teacher-forcing, as shown in Figure 5. (2) Based on our CCT, we present the Auto-regressive Policy (ARP), a model that learns to generate heterogeneous action sequences autoregressively for solving robotic manipulation tasks. The ARP architecture is summarized in Figure 3. (3) We evaluate the same ARP architecture across Push-T (Chi et al., 2023), ALOHA (Zhao et al., 2023), and RLBench (James et al., 2020), three environments with diverse manipulation tasks and control modalities, as outlined in Figure 2. Our study shows that ARP outperforms all environment-specific SoTAs, while being more efficient computationally and using smaller parameter sizes, as in Figure 6. In addition, we evaluate ARP with a real robot on a challenging, contact-rich nut-tightening task, as shown in Figure 9.
@@ 2 Related Work
Learning robotic manipulation from demonstrations. Imitation learning enables robots to learn to perform tasks demonstrated by experts (Zare et al., 2024). Recently, various methods have been developed for manipulation learning with different task constraints and control modalities. Notably, Chi et al. (2023) proposed the Diffusion Policy (DP) method for solving the Push-T task. Zhao et al. (2023) proposed the Action Chunking Transformer (ACT) for bi-manual manipulation tasks in the ALOHA environment. Goyal et al. (2024) proposed RVT-2 for language-conditioned tasks in the RLBench environment (James et al., 2020). We outline these environments and the corresponding state-of-the-art (SoTA) solutions in Figure 2 and Figure 7, respectively. In contrast, our proposed autoregressive policy is a universal architecture that outperforms each environment-specific SoTA on Push-T, ALOHA, and RLBench.
Autoregressive models for control tasks. Besides the pioneering Decision Transformer and Trajectory Transformer, recent works such as VIMA (Jiang et al., 2022), Gato (Reed et al., 2022), GR1 (Wu et al., 2023), OpenVLA (Kim et al., 2024) and ManipLLM (Li et al., 2024) have looked into designing autoregressive models for robotic tasks. Despite the impressive results, these approaches are limited to low-frequency control tasks that rely on end-effector waypoints (Kim et al., 2024). GR1 and OpenVLA only use autoregression in their LLM backbones and do not apply autoregression to solve control tasks. Most of the existing works require fine-tuning a large language model (LLM) such as LLaMA (Li et al., 2024) to include target end-effector poses within text-based responses or predict poses from LLM’s hidden features. The reliance on resource-intensive LLMs leads to large computational overhead, even for tasks that could be addressed with lightweight models. Without these constraints, our autoregressive policy outperforms SoTAs in multiple environments while being more efficient in computation and parameter sizes.
Hierarchical policies. Planning actions on multiple levels of abstraction is an important ability (Pateria et al., 2021). Existing methods generally separate the designs of low-level and high-level policies, and uses different modules for the different levels of abstraction (Le et al., 2018; Pateria et al., 2021; Belkhale et al., 2023; 2024; Chen et al., 2024b). This complicated procedure prohibits a wider application. In contrast, our autoregressive policy is able to capture the underlying causal dependencies in robotic tasks by predicting a sequence of actions of different levels of abstraction by using a single model, as shown by our action sequence designs for diverse environments in Figure 4.
Figure 3: Autoregressive Policy Architecture. A sequence of past actions and a chunk of empty tokens are concatenated and projected into embeddings. The empty tokens correspond to future actions, which are unknown and need to be predicted. These embeddings are fed into our Chunking Causal Transformer (CCT) along with the vision features of the current observation. CCT alternates between self-attention within the input embeddings and cross-attention with the vision features. Self-attention is causal for the input actions and bidirectional among the empty tokens. Distributions of future actions are decoded from the updated embeddings of the empty tokens.
@@ 3 Method
Action sequence modeling. Unlike natural language, robot actions lack a universal vocabulary. As shown in Figure 2, different robot tasks may require drastically different types of actions. Therefore, we represent actions as structured sequences whose formats are tailored for each family of tasks. Figure 4 showcases the formats of the action sequences generated in our real robot experiment, Push-T, ALOHA, and RLBench tasks.
Embedding and decoding heterogeneous actions. Language models map each word to a continuous vector called word embedding. The word embeddings of the input sentences are fed into a causal transformer. The distribution of the next word is decoded from the output embedding of the last word with a linear layer. Figure A2 and A3 illustrate our embedding and decoding methods for robot actions. Discrete actions are embedded by a table lookup on a weight matrix and decoded into a categorical distribution with a linear layer, similar to words in language modeling. Continuous actions are embedded with a linear layer and decoded into the parameters of a Gaussian mixture distribution with another linear layer. Actions that are defined as pixel coordinates are embedded by retrieving the pointwise features at the coordinates on a visual feature map. The output spatial distribution is obtained by multiplying the output embedding with the visual feature map, and converting the result into a 2d heatmap with the upsampling operator from RAFT (Teed & Deng, 2020).
@@
Chunking causal transformer. Figure A1 illustrates the essential difference between a causal transformer and our CCT. A causal transformer modifies the token embedding with causal attention so that the last token becomes the next token. Our CCT modifies the token embedding with causal attention for the action tokens $a_{i}$ (past actions) and bidirectional attention for the empty tokens $e_{i}$ (future actions). The empty tokens become the next tokens. This allows the prediction of a chunk of variable number of next tokens at once in a single forward pass by adding empty tokens, which enables action chunking during autoregressive generation. We study the impacts of our chunking autoregression in detail in Section 4. In ARP, CCT alternates between self-attention within the input embeddings and cross-attention with vision features, as in Figure 3. We extract vision features from a standard backbone identical to the ones used in SoTA methods, as detailed in section 4.
@@
To resolve this, we first divide the attention into three categories: (1) causal attention within the action tokens, (2) bidirectional attention within the empty tokens, and (3) causal attention between the empty and the action tokens, as marked in Figure 5. Figure A4 shows that the causal attention within action tokens is computed repeatedly. We avoid this redundancy by precomputing the causal attentions for all action tokens and caching the results. For the empty tokens, we combine their inner bidirectional attention and the causal attention toward cached action tokens. This enables a single forward pass of all tokens in three attention operations, regardless of the number of tokens. We name this procedure attention interleaving. Figure A4 demonstrates the reduced MACs of training with attention interleaving. We implement attention interleaving as an internal acceleration mechanism of the transformer layer, which is transparent to other network modules. Note that attention interleaving is only used during training and incurs no additional inference cost.
Inference. During the test rollouts, we extract vision tokens from the current observation and provide them as input to ARP, which then generates actions autoregressively by sampling from the decoded action distribution and appending the selected actions to the existing action sequence. This process of generating and executing actions is repeated until episode termination (success, failure, or reaching the time limit). Actions are generated according to the sequence formats shown in Figure 4. We manually set the chunk size for each type of action, and total sequence length for each task. We provide more implementation details and hyper-parameter values in Section 4 and Appendix A.2.
Figure 5: Training Chunking Causal Transformer (CCT) with Teacher-forcing. Causal transformers are trained efficiently with only a single forward pass for all tokens in a given sequence. However, suppose $a_{2},a_{3}$ and $a_{4}$ are in separate chunks, the CCT forward passes of predicting $a_{2},a_{3}$ and $a_{4}$ cannot be merged directly. Naively running separate passes significantly increases computation costs, as in Figure A4. With the proposed attention interleaving, we precompute and cache the causal attention results for all action tokens $a_{1},a_{2},a_{3}$. For the empty tokens $e_{2},e_{3},e_{4}$, we combine their bidirectional attention (lightgray) and the causal attention towards action tokens (green, blue). Since all the computations related to the action tokens are cached, we can update all the empty tokens in a single pass, regardless of the number of tokens. An example of attention interleaving is provided in the Video/attention-interleaving-tour.mp4 in the supplementary.
@@ 4.1 Comparison with State-of-the-Art
Setup. We compare the autoregressive policy (ARP) against the SoTA solutions in Push-T, ALOHA, and RLBench environments. Push-T is a single task. ALOHA consists of two tasks: insertion and cube transfer. RLBench includes 18 tasks, each with multiple language variants. These environments are illustrated in Figure 2 and Figure A5. For Push-T and ALOHA, we train a separate policy for each task. For RLBench, a single policy is trained for all 18 tasks. In Push-T, the policy observes the last two $96\times 96$ RGB frames, and predicts a window of future 2-d pointer positions as actions. In ALOHA, the policy observes the current $480\times 640$ RGB frame and predicts a window of future 14-dimensional joint positions. In RLBench, the policy observes four RGBD $128\times 128$ images and predicts the next target end-effector pose and gripper states. Existing SoTA techniques in these environments are outlined in Figure 7. We use the same vision backbones as the SoTA solutions to extract vision tokens, namely ResNet50 (He et al., 2016) for Push-T and ALOHA, and Multi-View Transformer (Goyal et al., 2023) for RLBench. We use the same training data, number of episodes, optimizer configuration, and evaluation frequency as the SoTA solutions. We detail the full list of hyper-parameters, such as the number of layers, sequence lengths, chunk sizes, and optimizer setups in Appendix A.2. Success rates for Push-T and RLBench are averaged over three independent runs. ALOHA’s results are averaged over five runs.
Results. Figure 6 shows that our autoregressive policy (ARP) outperforms environment-specific SoTAs while being more computationally efficient. Table A4 compares the per-task success rates of our ARP and RVT-2 (Goyal et al., 2024). In addition, we report the result of ARP+, which shares the same network definition with ARP but has more layers. The MACs / parameter sizes of RVT-2, ARP, ARP+ are 72.1M/57.1G, 71.9M/56.2G, and 73.8M/57.4G, respectively. Note that the originally reported success rate of RVT-2 was 81.6%. This was achieved by using a logical time-step as an input. Many existing works have followed this convention (Shridhar et al., 2023; Goyal et al., 2023). We train all our RLBench models without the input time-step. Thus, our RLBench models are more robust to variations of task initialization in real applications.